The force required to stretch a Hooke’s-law
spring varies from 0 N to 42.5 N as we stretch
the spring by moving one end 7.23 cm from
its unstressed position.
Find the force constant of the spring.
Answer in units of N/m.
Don't round answer
1
Expert's answer
2019-11-06T08:53:52-0500

The force constant of the spring:


"k=\frac{F}{x}"

"k=\frac{42.5}{0.0723}=588\frac{N}{m}"